### v1

@@ A U-statistic is an unbiased estimator for the parameter
$U^{(d)}_{\mathbf{X}_{I^{(d)}}}(k^{(d)})=\binom{n}{r^{\max}}^{-1}\sum_{1\leq t_{1}<\ldots<t_{r^{\max}}\leq n}k^{(d)}\left(\mathbf{X}_{t_{1}},\ldots,\mathbf{X}_{t_{r^{(d)}}}\right).$
Here, each of the summands in the upper representation shows up $\binom{n}{r^{\max}}/\binom{n}{r^{(d)}}$ times, but is then divided by this factor to ensure equality. Using this, $\widetilde{U}$ can be rewritten as
$\widetilde{U}=\binom{n}{r^{\max}}^{-1}\sum_{1\leq t_{1}<\ldots<t_{r^{\max}}\leq n}\left(\lambda_{1}k^{(1)}(\mathbf{X}_{I^{(1)},t_{1}},\ldots,\mathbf{X}_{I^{(1)},t_{r^{(1)}}})+\ldots+\lambda_{D}k^{(D)}(\mathbf{X}_{I^{(D)},t_{1}},\ldots,\mathbf{X}_{I^{(D)},t_{r^{(D)}}})\right)$
and is thus itself a U-statistic with a bounded and symmetric kernel. Hence, we can apply the univariate central limit theorem from Proposition A.2 to $\widetilde{U}$. Invoking the Cramér-Wold device (vandervaart2000, Chapter 2) completes the proof. ∎
@@ Appendix B Proofs and Additional Lemmas
For (iv), the expectation of $G_{X}(X)$ is calculated in neslehova2007, and the variance of $G_{X}(X)$ follows from neslehova2007 after correcting a typo (the square occurring in (27) has to be within the expectation) and noting that $\mathrm{E}[(F_{X}(X)-F_{X}(X^{-}))^{2}]=\mathrm{P}(X=X^{\prime}=X^{\prime\prime})$. ∎
@@
By the delta method (vandervaart2000, Theorem 3.1), we thus have
@@
By Theorem 3 in doukhan_mixing_1994, equation (1.11) in bradley2005 and the fact that mixing coefficients can only decrease when applied to functions of some original time series, it holds that
@@
The speed of convergence follows from the speed in which the empirical probability measure converges to the theoretical probability measure, also under Assumption 2 (doukhan1995invariance). Indexing
@@
where either $l=\rho$ or $m=\rho$ or both. The first term is $\mathcal{O}_{p}(1/\sqrt{n})$ because of the same arguments as for (B). The second term is also $\mathcal{O}_{p}(1/\sqrt{n})$ because the function class $\{y\mapsto G_{X,Y}(x,y),x\in\mathbb{R}\}$ satisfies the conditions in doukhan1995invariance and the empirical probability measure converges with $\sqrt{n}$ rate.
@@ Appendix C Cross-dependence of Bivariate Geometric Distibution
In order to illustrate our dependence concepts and the associated computations in the case of discrete random variables, let us consider the three-parameter bivariate geometric distribution for $(X,Y)$ as proposed by barbiero19, which we denote as $\mathrm{BGeom}(\pi_{X},\pi_{Y},\alpha)$ with parameters $\pi_{X},\pi_{Y}\in(0,1)$ and $\alpha\in[-1,1]$. The BGeom distribution is constructed by applying the Farlie–Gumbel–Morgenstern (FGM) copula (see nelsen06, p. 77) to the univariate geometric distributions $X\sim\mathrm{Geom}(\pi_{X})$ and $Y\sim\mathrm{Geom}(\pi_{Y})$, where $\alpha$ controls the extent of cross-dependence (with independence for $\alpha=0$). Note that the FGM copula can only model relatively weak dependence (see again nelsen06, p. 77), and in particular does not include the copulas representing perfect positive and negative dependence (the co- and countermonotonicity copulas) as limiting cases. Thus, this family of copulas and with it the BGeom distribution are of limited use for modeling. However, they are perfectly suited to our purpose since they allow for a closed-form computation of the formulas we want to illustrate. Note that this is also the reason why the dependence measures calculated below can only take values close to zero even in the continuous case and if $\alpha=\pm 1$. The CDF and PMF of the BGeom distribution are given by
@@
If the FGM-copula with dependence parameter $\alpha\in[-1,1]$ is applied to continuous $X,Y$, then it is known that $\tau(X,Y)=\frac{2}{9}\,\alpha$ and $\rho(X,Y)=\frac{3}{2}\,\tau(X,Y)=\frac{1}{3}\,\alpha$, see nelsen06. In the discrete case $\mathrm{BGeom}(\pi_{X},\pi_{Y},\alpha)$, however, we obtain another result, namely
@@ D.1.1 Considered DGPs
where $\circ$ denotes the binomial thinning operator (steutel79) and $\odot$ the signed binomial thinning operator (kim08). These thinning operators serve as integer-valued substitutes of multiplication (weiss2008c; weiss18). While the Poisson DGP (D.5) can be interpreted as the “normal case” in the count-data world, the Zipf DGP (D.6) causes heavy tails and non-existing moments (johnson05, Section 11.2.20), whereas the Skellam DGP (D.7) has $\mathbb{Z}$-valued outcomes (johnson05, Section 4.12.3) and allows to capture negative cross-dependence.
@@ D.1.2 Analysis of Simulation Results
where $\widehat{r}$ denotes the empirical Pearson correlation (serfling80, pp. 125–126).
@@
Next, we investigate the performance of 90% confidence intervals (CIs) for the different DGPs, where we use our proposed variance estimator from (16). While we considered all rank correlations for the task of testing for dependence in our simulation study, we restrict our analysis of CIs to such rank correlations that do not suffer (much) from attainability problems. For continuous DGPs, we simply use Kendall’s $\tau$ from Definitions 2.2 and 3.1 and Spearman’s $\rho$ from Definitions 2.3 and 3.2 again. For discrete DGPs, by contrast, recall our discussion in Sections 1 and 2, it is known that only Goodman-Kruskal’s $\gamma$ from Definitions 2.4 and 3.3 avoids any attainability problems (pohle2025), and the grade correlation $\rho_{b}$ from Definitions 2.6 and 3.5 at least mitigates them (neslehova2007).
@@ D.2.1 Considered DGPs
as counterparts to (D.1)–(D.3), and a bivariate version of the TEAR(1)(1) (transposed exponential AR) process of lawrance81
@@
As discrete AR(1)(1)-like processes, see weiss2008c; weiss18 for surveys, we use $(X_{i},Y_{i}):=\big(X_{i},\ \alpha\circ X_{i}+(1-\alpha)\circ u_{i}\big)$, $\alpha\in(0,1)$, together with the INAR(1)(1) (integer-valued AR) processes (kenzie85)
@@
In order to allow for negative dependence, we again replace the binomial thinning operator by its signed counterpart $(X_{i},Y_{i}):=\big(X_{i},\ \alpha\odot X_{i}+(\mathds{1}_{\{\alpha\geq 0\}}-\mathds{1}_{\{\alpha<0\}})\cdot(1-|\alpha|)\odot u_{i}\big)$, $\alpha\in(-1,1)$, with signed INAR(1)(1) (INARS(1)) processes (kim08)
@@ D.2.2 Analysis of Simulation Results
with $\sigma_{r,\textup{ind}}^{2}=\sum_{h=-\infty}^{\infty}r_{X}(h)r_{Y}(h)$, where $r_{X}(h)$ denotes the autocorrelation function (ACF) of $X$ at lag $h$ (haugh1976). We employ this limiting distribution and estimate the variance $\sigma_{r,\textup{ind}}^{2}$ with a HAC-type estimator with the same kernel and bandwidth choice as for our other HAC-type estimators.

### v2

@@ A U-statistic is an unbiased estimator for the parameter
$U^{(d)}_{\mathbf{X}_{I^{(d)}}}(k^{(d)})=\binom{n}{r^{\max}}^{-1}\sum_{1\leq i_{1}<\ldots<i_{r^{\max}}\leq n}\widetilde{k}^{(d)}\left(\mathbf{X}_{I^{(d)},i_{1}},\ldots,\mathbf{X}_{I^{(d)},i_{r^{max}}}\right),$
where
$\widetilde{k}^{(d)}\left(\mathbf{X}_{I^{(d)},i_{1}},\ldots,\mathbf{X}_{I^{(d)},i_{r^{max}}}\right):=\binom{r^{max}}{r^{(d)}}^{-1}\sum_{1\leq j_{1}<...<j_{r^{(d)}}\leq r^{max}}k^{(d)}\left(\mathbf{X}_{I^{(d)},i_{1}},\ldots,\mathbf{X}_{I^{(d)},i_{r^{(d)}}}\right).$
Here, each of the summands in the upper representation shows up $\binom{n}{r^{\max}}\binom{r^{\max}}{r^{(d)}}/\binom{n}{r^{(d)}}$ times, but is then divided by this factor to ensure equality. Using this, $\widetilde{U}$ can be rewritten as
$\widetilde{U}=\binom{n}{r^{\max}}^{-1}\sum_{1\leq i_{1}<\ldots<i_{r^{\max}}\leq n}\left(\lambda_{1}\widetilde{k}^{(1)}(\mathbf{X}_{I^{(1)},i_{1}},\ldots,\mathbf{X}_{I^{(1)},i_{r^{max}}})+\ldots+\lambda_{D}\widetilde{k}^{(D)}(\mathbf{X}_{I^{(D)},i_{1}},\ldots,\mathbf{X}_{I^{(D)},i_{r^{max}}})\right)$
and is thus itself a U-statistic with a bounded and symmetric kernel. Hence, we can apply the univariate central limit theorem from Proposition A.2 to $\widetilde{U}$. Invoking the Cramér-Wold device (Van der Vaart, 2000, Chapter 2) completes the proof. ∎
@@ Appendix B Proofs and Additional Lemmas
For (iv), the expectation of $G_{X}(X)$ is calculated in Nešlehová (2007, Equation (13)), and the variance of $G_{X}(X)$ follows from Nešlehová (2007, (27)) after correcting a typo (the square occurring in (27) has to be within the expectation) and noting that $\mathrm{E}[(F_{X}(X)-F_{X}(X^{-}))^{2}]=\mathrm{P}(X=X^{\prime}=X^{\prime\prime})$. ∎
@@
By the delta method (Van der Vaart, 2000, Theorem 3.1), we thus have
@@
By Theorem 3 in Doukhan (1994), equation (1.11) in Bradley (2005) and the fact that mixing coefficients can only decrease when applied to functions of some original time series, it holds that
@@
The speed of convergence follows from the speed in which the empirical probability measure converges to the theoretical probability measure, also under Assumption 2 (Doukhan et al., 1995). Indexing
@@
where either $l=\rho$ or $m=\rho$ or both. The first term is $\mathcal{O}_{p}(1/\sqrt{n})$ because of the same arguments as for (B). The second term is also $\mathcal{O}_{p}(1/\sqrt{n})$ because the function class $\{y\mapsto G_{X,Y}(x,y),x\in\mathbb{R}\}$ satisfies the conditions in Doukhan et al. (1995) and the empirical probability measure converges with $\sqrt{n}$ rate.
@@ Appendix C Cross-dependence of Bivariate Geometric Distibution
In order to illustrate our dependence concepts and the associated computations in the case of discrete random variables, let us consider the three-parameter bivariate geometric distribution for $(X,Y)$ as proposed by Barbiero (2019), which we denote as $\mathrm{BGeom}(\pi_{X},\pi_{Y},\alpha)$ with parameters $\pi_{X},\pi_{Y}\in(0,1)$ and $\alpha\in[-1,1]$. The BGeom distribution is constructed by applying the Farlie–Gumbel–Morgenstern (FGM) copula (see Nelsen, 2006, p. 77) to the univariate geometric distributions $X\sim\mathrm{Geom}(\pi_{X})$ and $Y\sim\mathrm{Geom}(\pi_{Y})$, where $\alpha$ controls the extent of cross-dependence (with independence for $\alpha=0$). Note that the FGM copula can only model relatively weak dependence (see again Nelsen, 2006, p. 77), and in particular does not include the copulas representing perfect positive and negative dependence (the co- and countermonotonicity copulas) as limiting cases. Thus, this family of copulas and with it the BGeom distribution are of limited use for modeling. However, they are perfectly suited to our purpose since they allow for a closed-form computation of the formulas we want to illustrate. Note that this is also the reason why the dependence measures calculated below can only take values close to zero even in the continuous case and if $\alpha=\pm 1$. The CDF and PMF of the BGeom distribution are given by
@@
If the FGM-copula with dependence parameter $\alpha\in[-1,1]$ is applied to continuous $X,Y$, then it is known that $\tau(X,Y)=\frac{2}{9}\,\alpha$ and $\rho(X,Y)=\frac{3}{2}\,\tau(X,Y)=\frac{1}{3}\,\alpha$, see Nelsen (2006, pp. 162, 174). In the discrete case $\mathrm{BGeom}(\pi_{X},\pi_{Y},\alpha)$, however, we obtain another result, namely
@@ D.1.1 Considered DGPs
where $\circ$ denotes the binomial thinning operator (Steutel and van Harn, 1979) and $\odot$ the signed binomial thinning operator (Kim and Park, 2008). These thinning operators serve as integer-valued substitutes of multiplication (Weiß, 2008, 2018). While the Poisson DGP (D.5) can be interpreted as the “normal case” in the count-data world, the Zipf DGP (D.6) causes heavy tails and non-existing moments (Johnson et al., 2005, Section 11.2.20), whereas the Skellam DGP (D.7) has $\mathbb{Z}$-valued outcomes (Johnson et al., 2005, Section 4.12.3) and allows to capture negative cross-dependence.
@@ D.1.2 Analysis of Simulation Results
where $\widehat{r}$ denotes the empirical Pearson correlation (Serfling, 1980, pp. 125–126).
@@
Next, we investigate the performance of 90% confidence intervals (CIs) for the different DGPs, where we use our proposed variance estimator from (16). While we considered all rank correlations for the task of testing for dependence in our simulation study, we restrict our analysis of CIs to such rank correlations that do not suffer (much) from attainability problems. For continuous DGPs, we simply use Kendall’s $\tau$ from Definitions 2.2 and 3.1 and Spearman’s $\rho$ from Definitions 2.3 and 3.2 again. For discrete DGPs, by contrast, recall our discussion in Sections 1 and 2, it is known that only Goodman-Kruskal’s $\gamma$ from Definitions 2.4 and 3.3 avoids any attainability problems (Pohle and Wermuth, 2025), and the grade correlation $\rho_{b}$ from Definitions 2.6 and 3.5 at least mitigates them (Nešlehová, 2007).
@@ D.2.1 Considered DGPs
as counterparts to (D.1)–(D.3), and a bivariate version of the TEAR(1)(1) (transposed exponential AR) process of Lawrance and Lewis (1981)
@@
As discrete AR(1)(1)-like processes, see Weiß (2008, 2018) for surveys, we use $(X_{i},Y_{i}):=\big(X_{i},\ \alpha\circ X_{i}+(1-\alpha)\circ u_{i}\big)$, $\alpha\in(0,1)$, together with the INAR(1)(1) (integer-valued AR) processes (McKenzie, 1985)
@@
In order to allow for negative dependence, we again replace the binomial thinning operator by its signed counterpart $(X_{i},Y_{i}):=\big(X_{i},\ \alpha\odot X_{i}+(\mathds{1}_{\{\alpha\geq 0\}}-\mathds{1}_{\{\alpha<0\}})\cdot(1-|\alpha|)\odot u_{i}\big)$, $\alpha\in(-1,1)$, with signed INAR(1)(1) (INARS(1)) processes (Kim and Park, 2008)
@@ D.2.2 Analysis of Simulation Results
with $\sigma_{r,\textup{ind}}^{2}=\sum_{h=-\infty}^{\infty}r_{X}(h)r_{Y}(h)$, where $r_{X}(h)$ denotes the autocorrelation function (ACF) of $X$ at lag $h$ (Haugh, 1976). We employ this limiting distribution and estimate the variance $\sigma_{r,\textup{ind}}^{2}$ with a HAC-type estimator with the same kernel and bandwidth choice as for our other HAC-type estimators.